ET= 24 v E- 1;= R = 22 IT- RT = P= E]= I= R = 122 Ey= I;- R = 42
Answer #1

R2 and R3 are in series

R = R2 + R3 = 2 + 4 = 6 ohm

12 and 6 ohm are in parallel

R(equi) = 12*6 / (12 + 6) = 4 ohm

SO Rt = 4 ohm

It = Et/Rt = 24 / 4 = 6 A

Pt = Et*It = 24*6 = 144 W

E1 = 24 V

I1 = E1/R1 = 24/12 = 2A

I2 = I3 = I - I1 = 6 - 2 = 4A

E2 = I2*R2 = 4*2 = 8 V

E3 = I3*R3 = 4*4 = 16 V